.65 mL of .24M NaOH are added to 21.6 mL of buffer calculate pH of buffer prior to base addition, after and pH change.
what if instead of buffer it is 21.6 mL of distilled water. ph before, after and the change. 3 sig figs.

buffer was made from .196M HCHOO .271M NaCHOO. with a given ph of 4.4
Answer #1

Given Buffer pH = 4.4 Chcool] - 0.196 M (Na CHOD] = 0.271 M pH of buffer is given by Hendersons Equation pH = pka + log [NapH before addition de base - 4.40 pka of acid = 4.26 Naon pH after addition Volume of of 0.65 mL 0.24 M Naol = 0.65m milimolepH = pka + log CHCOONa] Cacoon] é 4.26 + log 6.0096 4.0776 = 4.26 + 0.168 _ After addition of Naon) pH = 4.428 m 4.43 ph afteNo. o milimoles do Naon - Molavity x volume = 0.247 x 0.65 m2 - 0.156 milimoles. olarity of NaOH = milimoles of Naoh Total vo11.85 pH after addition of Naon change in pH = 11.85-7 = 4.85 change in pH = 4.85